# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ B.3 Logistic-beta dependent Dirichlet process and related models
Consider the single-atoms logistic-beta dependent Dirichlet process, so that atoms $\theta_{h}$ are i.i.d. samples from the base measure $G_{0}$. Let $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$ and $F_{B}$ be the c.d.f. of $\mathrm{Beta}(1,b)$. Consider an $n$-dimensional continuous random variable $U=(F_{B}[\sigma\{\eta(x_{1})\}],\dots,F_{B}[\sigma\{\eta(x_{n})\}])$ where its cumulative distribution $C_{x_{1},\dots,x_{n}}$ is a copula since $U$ has uniform marginals by construction. It is clear that $U$ has a positive density w.r.t. Lebesgue measure since $F_{B}$ and $\sigma$ are all continuous and injective.
Now consider a collection of copula functions $\mathcal{C}=\{C_{x_{1},\dots,x_{n}}:x_{1},\dots,x_{n}\in\mathscr{X},n>1\}$. Since $\sigma$ is the c.d.f. of a standard logistic random variable, the collection of copula functions satisfies conditions of Corollary 1 of Barrientos et al., (2012). Also, the stick-breaking ratios of logistic-beta dependent Dirichlet process $\sigma\{\eta_{h}(x)\}$ are determined by the set of copulas $\mathcal{C}$ and the set of beta marginal distributions with parameters $1$ and $b$. Therefore, the full weak support property of the logistic-beta dependent Dirichlet process follows from Theorem 3 of Barrientos et al., (2012). For a general logistic-beta dependent Dirichlet process where atoms and weights both depend on $x$, the full weak support condition is satisfied as long as the atom process can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure.
Appendix C Additional posterior computation strategies
C.1 Correlation kernels with low-rank structures
Thanks to the normal variance-mean mixture construction of the logistic-beta process, many existing scalable Gaussian process methods that preserve marginal variances can be easily applied to the logistic-beta process. When the dimension of normalized feature map $q$ is less than data size $n$, adopting normalized feature map kernel offers dimension reduction using basis functions (Cressie et al.,, 2022), which is apparent from two representations (6)–(7) in Proposition 3. Specifically, denoting $\Phi=\{\phi_{k}(x_{i})\}_{ik}$ be an $n\times q$ basis matrix with $i$th row $\phi(x_{i})$, the Step 3 of Algorithm 9 becomes sampling $\gamma$ from $q$-dimensional multivariate normal instead of $n$-dimensional multivariate normal,
$\gamma\mid-\sim N_{q}\left((I_{q}+\lambda\Phi^{\top}\Omega\Phi)^{-1}\lambda^{1/2}\Phi^{\top}\{(z_{1:n}-0.51_{n})-0.5\lambda(a-b)\omega\},(I_{q}+\lambda\Phi^{\top}\Omega\Phi)^{-1}\right)$
and set $\eta=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma$ using representation (7).
@@ C.1 Correlation kernels with low-rank structures
Another example of correlation kernel with low-rank structures is the modified predictive process (Finley et al.,, 2009). From a parent kernel $\mathcal{R}$ with $q$ knot locations $u_{1},\dots,u_{q}\in\mathscr{X}$, a new correlation kernel $\tilde{\mathcal{R}}$ with a low-rank structure can be defined as
~$\tilde{\mathcal{R}}(x,x^{\prime})=r(x)^{\mathrm{\scriptscriptstyle T}}R_{uu}^{-1}r(x^{\prime})+\mathds{1}(x=x^{\prime})(1-r(x)^{\mathrm{\scriptscriptstyle T}}R_{uu}^{-1}r(x^{\prime}))$ (14)
@@
where $r(x)=[\mathcal{R}(x,u_{k})]_{k=1}^{q}$ is a length $q$ vector as a function of $x$ and $R_{uu}=[\mathcal{R}(u_{k},u_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a fixed $q\times q$ knot correlation matrix; see also Quiñonero-Candela and Rasmussen, (2005). Thanks to the Woodbury matrix identity and matrix determinant lemma, posterior computation with a low-rank correlation kernel can be carried out in a highly efficient manner. This includes fast evaluation of $n$-dimensional multivariate normal density in Step 2 of Algorithm 9, as the covariance matrix has a “low-rank plus diagonal” structure. In addition to the modified predictive process (14), other options include full-scale approximation (Sang and Huang,, 2012) or multi-resolution approximation (Katzfuss,, 2017) among many others.
@@
C.2 Particle Gibbs sampler
@@
In Step 2 of Algorithm 9, it is possible to use a particle Gibbs sampler (Andrieu et al.,, 2010) to potentially improve the mixing behaviour of $\lambda$. The adaptive Pólya proposal scheme can be used to choose the particle proposal. In Step 2 of Algorithm 9, instead of drawing a single candidate, the particle Gibbs sampler first draws $N$ candidates $\lambda^{(1)},\dots,\lambda^{(N)}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$ and sets the current state as the last candidate $\lambda^{(N+1)}=\lambda$, where $(a^{\prime},b^{\prime})$ are chosen based on adaptive Pólya proposal scheme. Then, unnormalized importance weights $w_{k}=\exp\{-\lambda^{(k)}(ab-a^{\prime}b^{\prime})/2\}g(\lambda^{(k)})$ for $k=1,\dots,N+1$ are calculated, and $\lambda$ is sampled among $N+1$ candidates according to probabilities proportional to $w_{k}$’s. This particle Gibbs sampler also avoids the evaluation of the Pólya density and potentially leads to better mixing with large $N$ but at an increased cost compared to Metropolis-Hastings.
C.3 Challenges for developing rejection samplers involving Pólya densities
@@
Appendix D Details of simulation analysis and application study
D.1 Simulation studies
@@
Table D.1: Simulation results under various latent logistic-beta process inference algorithm settings with latent logistic-beta process data generation scenarios. Entries shown are average over 100 replicates, with Monte Carlo standard error shown in parentheses. LBP, logistic-beta process; M-H, Metropolis-Hastings; ESS, effective sample size of $\lambda$; ESS/sec, ESS divided by running time in seconds.
Data generation Algorithm (blocked) ESS ESS/sec Acceptance rate (%)
Latent LBP, $\varrho=0.1$ Indep. M-H Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
@@
Particle Gibbs Adapted 358.10 (18.57) 2.33 (0.12) 84.50 (0.42)
@@
Latent LBP, $\varrho=0.2$ Indep. M-H Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
@@
Particle Gibbs Adapted 360.48 (21.29) 1.96 (0.12) 86.11 (0.40)
@@
Latent LBP, $\varrho=0.4$ Indep. M-H Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
@@
Particle Gibbs Adapted 468.76 (25.98) 2.52 (0.14) 87.09 (0.29)

### v3

@@ Acknowledgements
    We first show Theorem 4.2 and specialize to Theorem 4.1 later. The proof is an extended version of the Theorem 1 of Dunson and Park, (2008). Denote $V_{h}=\sigma\{\eta_{h}(x)\}$ and $V_{h}^{\prime}=\sigma\{\eta_{h}(x^{\prime})\}$, which marginally follow $V_{h}\sim\mathrm{Beta}(1,b)$ and $V_{h}^{\prime}\sim\mathrm{Beta}(1,b)$. Since the dependent Dirichlet process prior implies $G_{x}$ marginally follows a Dirichlet process with concentration parameter $b$, we have $E\{G_{x}(B)\}=G^{0}(B)$ and $\operatorname*{var}\{G_{x}(B)\}=G^{0}(B)\{1-G^{0}(B)\}/(1+b)$ for any $x$ (Müller et al.,, 2015). In addition to the notation $\mu(x,x^{\prime})=E[\sigma\{\eta(x)\}\sigma\{\eta(x^{\prime})\}]$ where $\eta(x)\sim{\textsc{lbp}}(1,b,\mathcal{R})$, denote $\mu(x)=E[\sigma\{\eta(x)\}]$ which is simply $\mu(x)=1/(1+b)$ for any $x$. Recall the notation $G^{0}_{x,x^{\prime}}(B)=\operatorname*{pr}(\theta_{h}(x)\in B,\theta_{h}(x^{\prime})\in B)$, which does not depends on $h$ since atom processes are i.i.d. across $h$.
    The mixed moment is
    $E\{G_{x}(B)G_{x^{\prime}}(B)\}$
    =\,\,= $E\left[\sum_{h=1}^{\infty}V_{h}\left\{\prod_{l=1}^{h-1}(1-V_{l})\right\}\delta_{\theta_{h}(x)}(B)\times\sum_{h=1}^{\infty}V_{h}^{\prime}\left\{\prod_{l=1}^{h-1}(1-V_{l}^{\prime})\right\}\delta_{\theta_{h}(x^{\prime})}(B)\right]$
    =\,\,= $G_{x,x^{\prime}}^{0}(B)\underbrace{E\left[\sum_{h=1}^{\infty}V_{h}V_{h}^{\prime}\prod_{l=1}^{h-1}(1-V_{l})(1-V_{l}^{\prime})\right]}_{(*)}$
    $+G^{0}(B)^{2}\underbrace{E\left[\sum_{h\geq 1,k\geq 1,h\neq k}V_{h}V_{k}^{\prime}\prod_{l=1}^{h-1}(1-V_{l})\prod_{m=1}^{k-1}(1-V^{\prime}_{m})\right]}_{(**)}$
    where equality follows from atom processes and weight processes being independent. Since $(V_{h},V_{h}^{\prime})$ is independent of $(V_{k},V_{k}^{\prime})$ for any $h\neq k$, the term ()(*)( ∗ ) becomes
    ((*)( ∗ ) $=\sum_{h=1}^{\infty}E(V_{h}V_{h}^{\prime})\prod_{l=1}^{h-1}E\{(1-V_{l})(1-V_{l}^{\prime})\}$
    $=\sum_{h=1}^{\infty}\mu(x,x^{\prime})\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{h-1}$
    $=\frac{\mu(x,x^{\prime})}{\mu(x)+\mu(x^{\prime})-\mu(x,x^{\prime})}=\frac{(1+b)}{2/\mu(x,x^{\prime})-(1+b)}$
    The term ()(**)( ∗ ∗ ) becomes, since $(V_{h},V_{h}^{\prime})$ is independent of $(V_{k},V_{k}^{\prime})$ for any $h\neq k$,
    ((**)( ∗ ∗ ) $=E\sum_{h=1}^{\infty}\sum_{k=1}^{h-1}V_{h}\left[\prod_{m=1}^{k-1}(1-V_{m})(1-V_{m}^{\prime})\right](V_{k}^{\prime}-V_{k}V_{k}^{\prime})\prod_{s=k+1}^{h-1}(1-V_{s})$
    $+E\sum_{h=1}^{\infty}\sum_{k=h+1}^{\infty}V_{k}^{\prime}\left[\prod_{l=1}^{h-1}(1-V_{l})(1-V_{l}^{\prime})\right](V_{h}-V_{h}V_{h}^{\prime})\prod_{s=h+1}^{k-1}(1-V_{s}^{\prime})$
    $=\sum_{h=1}^{\infty}\sum_{k=1}^{h-1}\mu(x)\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{k-1}\{\mu(x^{\prime})-\mu(x,x^{\prime})\}\{1-\mu(x)\}^{h-k-1}$
    $+\sum_{h=1}^{\infty}\sum_{k=h+1}^{\infty}\mu(x^{\prime})\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{h-1}\{\mu(x)-\mu(x,x^{\prime})\}\{1-\mu(x^{\prime})\}^{k-h-1}$
    $=\sum_{k=1}^{\infty}\sum_{h=k+1}^{\infty}\mu(x)\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{k-1}\{\mu(x^{\prime})-\mu(x,x^{\prime})\}\{1-\mu(x)\}^{h-k-1}$
    $+\sum_{h=1}^{\infty}\sum_{k=h+1}^{\infty}\mu(x^{\prime})\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{h-1}\{\mu(x)-\mu(x,x^{\prime})\}\{1-\mu(x^{\prime})\}^{k-h-1}$
    $=\sum_{k=1}^{\infty}\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{k-1}\{\mu(x^{\prime})-\mu(x,x^{\prime})\}$
    $+\sum_{h=1}^{\infty}\{1-\mu(x)-\mu(x^{\prime})+\mu(x,x^{\prime})\}^{h-1}\{\mu(x)-\mu(x,x^{\prime})\}$
    $=\frac{\mu(x)+\mu(x^{\prime})-2\mu(x,x^{\prime})}{\mu(x)+\mu(x^{\prime})-\mu(x,x^{\prime})}$
    where reordering is justified as the series is absolutely convergent. Combining,
    $E\{G_{x}(B)G_{x^{\prime}}(B)\}=\frac{G_{x,x^{\prime}}^{0}(B)\mu(x,x^{\prime})}{\mu(x)+\mu(x^{\prime})-\mu(x,x^{\prime})}+G^{0}(B)^{2}\frac{\mu(x)+\mu(x^{\prime})-2\mu(x,x^{\prime})}{\mu(x)+\mu(x^{\prime})-\mu(x,x^{\prime})}$
    and thus, using $E\{G_{x}(B)\}=E\{G_{x^{\prime}}(B)\}=G^{0}(B)$ and $\mu(x)=\mu(x^{\prime})=1/(1+b)$,
    $\operatorname*{cov}\{G_{x}(B),G_{x^{\prime}}(B)\}=\frac{\{G_{x,x^{\prime}}^{0}(B)-G^{0}(B)^{2}\}\mu(x,x^{\prime})}{2/(1+b)-\mu(x,x^{\prime})}$
    and using $\operatorname*{var}\{G_{x}(B)\}=G^{0}(B)(1-G^{0}(B))/(1+b)$,
    $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}=\frac{G_{x,x^{\prime}}^{0}(B)-G^{0}(B)^{2}}{G^{0}(B)-G^{0}(B)^{2}}\frac{(1+b)^{2}}{2/\mu(x,x^{\prime})-(1+b)}$
    Now consider Theorem 4.1, the single-atoms DDP. We have $\theta_{h}(x)=\theta_{h}(x^{\prime})$ for any $x,x^{\prime}$ and $h$ and thus $G_{x,x^{\prime}}^{0}(B)=\operatorname*{pr}(\theta_{h}(x)\in B,\theta_{h}(x^{\prime})\in B)=G^{0}(B)$, so that first term cancels out and correlation does not depend on the evaluation set $B$, which gives
    $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}={(1+b)^{2}}/\left\{2\mu(x,x^{\prime})^{-1}-(1+b)\right\}.$
    Since atoms are common in $G_{x}$ and $G_{x^{\prime}}$, the tie probability is the same as the probability of picking the same label, corresponding to the term ()(*)( ∗ ) in the previous display. Thus, for $\vartheta\sim G_{x}$, $\vartheta^{\prime}\sim G_{x^{\prime}}$, we have
    $\operatorname*{pr}(\vartheta=\vartheta^{\prime}\mid G_{x},G_{x^{\prime}})={(1+b)}/\left\{2\mu(x,x^{\prime})^{-1}-(1+b)\right\}$
    which completes the proof. ∎
    S.1.3.3 Proof of Theorem 4.4
    Proof.
    If $\mathcal{R}(x,x^{\prime})$ is continuous at every point such that $x=x^{\prime}$, the covariance function $\operatorname*{cov}\{\eta(x),\eta(x^{\prime})\}$ is a linear function of $\mathcal{R}(x,x^{\prime})$ (Proposition 2.2) and thus continuous at every point $x=x^{\prime}$, i.e. the logistic-beta process is mean square continuous (Adler,, 1981, p.26). By Wise, (1977), its logistic transformation $V(\cdot)=\sigma\{\eta(\cdot)\}$ is also mean square continuous, which implies that $E\{V(x)V(x^{\prime})\}\to E\{V(x^{\prime})^{2}\}$ as $x\to x^{\prime}$. Since $V(\cdot)$ has $\mathrm{Beta}(1,b)$ marginal distribution, $E\{V(x)^{2}\}=2/\{(b+1)(b+2)\}$ for any $x\in\mathcal{X}$. Therefore, $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}\to(1+b)^{2}/\{(b+1)(b+2)-(1+b)\}=1$ as $x\to x^{\prime}$. ∎
    S.1.3.4 Details on related models
    Next, we present the prior correlation structure under other various dependent Dirichlet process models. In addition to the logistic-beta dependent Dirichlet process (M1), we provide a derivation of correlation lower bounds under three different dependent Dirichlet process models: (M2) squared AR(1) process-based dependent Dirichlet process (DeYoreo and Kottas,, 2018), (M3) time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012), and (M4) copula-based dependent Dirichlet process. Similar to the logistic-beta dependent Dirichlet process, these models induce dependence through a stochastic process on stick-breaking ratios that have beta marginals but with different constructions.
    First, from DeYoreo and Kottas, (2018, Appendix A.1.), the $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ under the squared ar(1) process-based dependent Dirichlet process has the greatest lower bound of $b^{1/2}(b+1)(b+2)^{1/2}-b(b+2)$ when autoregressive parameter converges to 0. This implies $E\{V(x)V(x^{\prime})\}$ has the infimum of $\{b^{3/2}+(1-b)(b+2)^{1/2}\}/\{(b+1)(b+2)^{1/2}\}$, using $E\{V(x)\}=1/(b+1)$ and $\operatorname*{var}\{V(x)\}=b/((b+1)^{2}(b+2))$ for any $x$ under beta distributed $V(x)$. This leads to the greatest lower bound of correlation between random probability measures $(1+b)/([2(b+2)^{1/2}/\{b^{3/2}+(1-b)(b+2)^{1/2}\}]-1)$, which converges to $0.75$ as $b\to\infty$.
    Next, the infimum of $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ under the time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012) is when $V(x)$ and $V(x^{\prime})$ are independent so that $\inf\operatorname*{corr}\{V(x),V(x^{\prime})\}=0$. Thus, $E\{V(x)V(x^{\prime})\}$ has the infimum of $1/(b+1)^{2}$, which leads to $\inf\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}=(1+b)/(1+2b)$. This fact is also noted in Nieto-Barajas et al., (2012) as a special case when stick-breaking ratios are independent, and the same fact also applies to the dependent Dirichlet process proposed by Taddy, (2010) when the autoregressive parameter converges to 0.
    Finally, we analyze a copula-based single-atoms dependent Dirichlet process, with beta stick-breaking ratios constructed through a copula that induces a stochastic process with $\mathrm{Beta}(1,b)$ marginals. Finding infimum of $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ is equivalent to finding the infimum of correlation between stick-breaking ratios $\operatorname*{corr}\{V(x),V(x^{\prime})\}$, which is achieved at the Fréchet lower bound copula (Embrechts et al.,, 2002). In other words, minimum correlation is attained when $V(x)$ and $V(x^{\prime})$ satisfy $V(x^{\prime})=F_{B}^{-1}(1-F_{B}(V(x)))$ (counter-monotonic), where $F_{B}$ is the c.d.f. of $\mathrm{Beta}(1,b)$.
    Although the analytic form of $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ under common atoms logistic-beta dependent Dirichlet process and copula-based dependent Dirichlet process are not readily available due to the nonlinear transformations involved, those can be estimated with Monte Carlo simulation where we used 5,000 simulated datasets of size 1,000 to calculate the correlation estimates.
    Figure S.1.1: Illustration of different correlation lower bounds under four different single-atoms dependent Dirichlet process. (Left) Infimum of correlation between random probability measures under models M1 (solid, black), M2 (dashed, red), M3 (dotted, purple), and M4 (dot-dashed, blue). (Right) Scatterplot of bivariate betas with $\mathrm{Beta}(1,2)$ marginals, corresponding to circles of the left panel.
    Figure S.1.1 illustrates different correlation lower bounds under four different single-atoms dependent Dirichlet processes with concentration parameter $b=2$. The difference in the induced dependence structure is clearly seen from scatterplots, where 1,000 random samples are drawn from a corresponding bivariate beta distribution that induces minimum correlation from models M1–M4. The Fréchet lower bound clearly shows counter-monotonicity between two beta random variables, which induces minimum possible correlation. Transformation of logistic-beta also can induce a negative correlation between beta random variables, while others cannot. In summary, logistic-beta dependent Dirichlet processes can capture a wide range of dependence when $b$ is small to moderate.
    In addition, Figure S.1.2 illustrates how various correlations change as a function of distance between $x$ and $x^{\prime}$ under Matérn correlation kernel with different smoothness parameters $\nu\in\{0.5,1.5,2.5\}$. This includes correlation between logistic-beta realizations $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}$, correlation between their logistic transformations $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ which are used as stick-breaking ratios, and correlation between random probability measures under single-atoms LB-DDP. In general, the correlation decay pattern of $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ and $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ are similar to the decay pattern of correlation between logistic-beta realizations $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}$, but in different scales depending on the choice of beta parameters $(a,b)$.
    Figure S.1.2: Correlation as a function of $d=\|x-x^{\prime}\|$ under Matérn kernel $\mathcal{R}_{M}(x,x^{\prime})=2^{1-\nu}\Gamma(\nu)^{-1}(d/\varrho)^{\nu}K_{\nu}(d/\varrho)$ with $(\varrho,\nu)\in\{(1/3,1/2),(4/19,3/2),(10/59,5/2)\}$. Each row corresponds to parameters $(a,b)\in\{(1,1/5),(1,1),(1,2),(1,10)\}$. (Left) $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}$, the correlation between logistic-beta process realizations. When $a=b$, we have $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}=\mathcal{R}_{M}(x,x^{\prime})$; see Proposition 2.2. (Center) $\operatorname*{corr}\{V(x),V(x^{\prime})\}$ with $V(x)=\sigma\{\eta(x)\}$, the correlation between stick-breaking ratios. (Right) $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$, the correlation between random probability measures under the single-atoms logistic-beta dependent Dirichlet process. Dot-dashed horizontal lines (gray) correspond to $\inf\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ when $\mathcal{R}(x,x^{\prime})=-1$. The range parameters $\varrho$ are chosen such that when $a=b$, $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}$ at distance 1 becomes approximately 0.05 (effective range is 1).
    Appendix S.2 Additional posterior computation strategies
    S.2.1 Correlation kernels with low-rank structures
@@
    Thanks to the normal variance-mean mixture construction of the logistic-beta process, many existing scalable Gaussian process methods that preserve marginal variances can be easily applied to the logistic-beta process. In addition to the correlation kernel constructed from normalized feature maps, another example of correlation kernel with low-rank structures is the modified predictive process (Finley et al.,, 2009). From a parent kernel $\mathcal{R}$ with $q$ knot locations $u_{1},\dots,u_{q}\in\mathscr{X}$, a new correlation kernel $\tilde{\mathcal{R}}$ with a low-rank structure can be defined as
    ~$\tilde{\mathcal{R}}(x,x^{\prime})=\mathbf{r}(x)^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{uu}^{-1}\mathbf{r}(x^{\prime})+\mathds{1}(x=x^{\prime})(1-\mathbf{r}(x)^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{uu}^{-1}\mathbf{r}(x^{\prime}))$ (S.1)
@@
    where $\mathbf{r}(x)=[\mathcal{R}(x,u_{k})]_{k=1}^{q}$ is a length $q$ vector as a function of $x$ and a matrix $\mathbf{R}_{uu}=[\mathcal{R}(u_{k},u_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a fixed $q\times q$ knot correlation matrix; see also Quiñonero-Candela and Rasmussen, (2005). Thanks to the Woodbury matrix identity and matrix determinant lemma, posterior computation with a low-rank correlation kernel can be carried out in a highly efficient manner. This includes fast evaluation of $n$-dimensional multivariate normal density in Step 2 of Algorithm 9, as the covariance matrix has a “low-rank plus diagonal” structure. In addition to the modified predictive process (S.1), other options include full-scale approximation (Sang and Huang,, 2012) or multi-resolution approximation (Katzfuss,, 2017) among many others.
@@
    S.2.2 Adaptive Pólya proposal and particle Gibbs sampler
    We describe the details of the adaptive Pólya proposal scheme, on choosing the proposal distribution parameters $(a^{\prime},b^{\prime})$ and sampling $\lambda$ from a distribution that involves Pólya prior density ${\pi_{\textsc{po}}}(\lambda;a,b)$, such as expression (9) in Step 2 of Algorithm 9. There are many ways to incorporate previous samples’ information in the choice of $(a^{\prime},b^{\prime})$. An example is the moment matching method with a running average of $\lambda$, denoted as $\bar{\lambda}$, that is recursively updated as $\bar{\lambda}\leftarrow\bar{\lambda}(m-1)/m+\lambda/m$ in $m$th iteration of the algorithm, leading to adaptive Metropolis-within-Gibbs (Roberts and Rosenthal,, 2009). Assuming $a^{\prime}\leq b^{\prime}$ without loss of generality due to the symmetry of the Pólya distribution in its arguments, which implies $a^{\prime}\leq c/2$, the $\mathrm{Polya}(a^{\prime},c-a^{\prime})$ has a mean $h(a^{\prime})=2\{\psi(a^{\prime})-\psi(c-a^{\prime})\}/(2a^{\prime}-c)$ (or $h(a^{\prime})=2\psi^{\prime}(c/2)$ if $a^{\prime}=c/2$) that is decreasing on $a^{\prime}\in(0,c/2]$. Thus, we set $a^{\prime}=(a+b)/2$ if $\bar{\lambda}$ is less than or equal to $2\psi^{\prime}((a+b)/2)$, and otherwise we set $a^{\prime}$ to the solution of $h(x)=\bar{\lambda}$ which can be easily obtained with numerical optimization algorithms.
@@
    In Step 2 of Algorithm 9, it is possible to use a particle Gibbs sampler (Andrieu et al.,, 2010) instead of Metropolis-Hastings to potentially improve the mixing behavior of $\lambda$. The adaptive Pólya proposal scheme can be used to choose the particle proposal. In Step 2 of Algorithm 9, instead of drawing a single candidate, the particle Gibbs sampler first draws $N$ candidates $\lambda^{(1)},\dots,\lambda^{(N)}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$ and sets the current state as the last candidate $\lambda^{(N+1)}=\lambda$, where $(a^{\prime},b^{\prime})$ are chosen based on adaptive Pólya proposal scheme. Then, unnormalized importance weights $w_{k}=\exp\{-\lambda^{(k)}(ab-a^{\prime}b^{\prime})/2\}\mathcal{L}(\lambda^{(k)})$ for $k=1,\dots,N+1$ are calculated, and $\lambda$ is sampled among $N+1$ candidates according to probabilities proportional to $w_{k}$’s. This particle Gibbs sampler also avoids the evaluation of the Pólya density and potentially leads to better mixing with large $N$ but at an increased cost compared to Metropolis-Hastings.
    S.2.3 Prior for beta parameters (a,b)
    We also consider a further hierarchical setting where beta parameters $(a,b)$ are no longer fixed but unknown. For simplicity, consider the nonparametric binary regression model where some proper prior $\pi(a,b)$ is placed on $(a,b)$. Although the inference could be carried out by having an additional step of sampling $(a,b)\mid\lambda$ in Algorithm 9 using Metropolis-Hastings, it involves the evaluation of Pólya densities which can cause numerical issues and can lead to slow mixing of the overall algorithm.
    We describe how inference can be still carried out without evaluation of Pólya densities and jeopardizing the overall mixing behavior. This is done by expanding the block sampling of $(\lambda,\bm{\eta})$ in Algorithm 9 to the block sampling of $(a,b,\lambda,\bm{\eta})$, by sequential updates from $\pi(a,b,\lambda\mid\bm{\omega},\mathbf{z})$, and $\pi(\bm{\eta}\mid a,b,\lambda,\bm{\omega},\mathbf{z})$. To update $(a,b,\lambda)$ jointly, one can use particle marginal Metropolis-Hastings algorithm (Andrieu et al.,, 2010). With $N$ number of particles, it proceeds as: (A) draw a new candidate $(a^{\star},b^{\star})\sim\tilde{q}(a^{\star},b^{\star}\mid a,b)$ with some proposal distribution $\tilde{q}$, (B) draw new particles $\lambda^{\star(1)},\dots,\lambda^{\star(N)}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Polya}(a^{\star},b^{\star})$, (C) Draw a candidate $\lambda^{\star}$ among $\lambda^{\star(1)},\dots,\lambda^{\star(N)}$ with probability proportional to $\mathcal{L}(\lambda^{\star(\ell)})$, $\ell=1,\dots,N$, and (D) Accept $(a^{\star},b^{\star},\lambda^{\star})$ and set of particles $\{\lambda^{\star(\ell)}\}_{\ell=1}^{N}$ with probability
    $\min\left\{1,\frac{\pi(a^{\star},b^{\star})\sum_{\ell=1}^{N}\mathcal{L}(\lambda^{\star(\ell)})}{\pi(a,b)\sum_{\ell=1}^{N}\mathcal{L}(\lambda^{(\ell)})}\times\frac{\tilde{q}(a,b\mid a^{\star},b^{\star})}{\tilde{q}(a^{\star},b^{\star}\mid a,b)}\right\},$
    otherwise keep $(a,b,\lambda)$ and current set of particles $\{\lambda^{(\ell)}\}_{\ell=1}^{N}$. The joint update of $(a,b,\lambda)$ from collapsed conditional $\pi(a,b,\lambda\mid\bm{\omega},\mathbf{z})$, where $\bm{\eta}$ is marginalized out, is made possible due to normal-normal conjugacy. Updating $\bm{\eta}$ is the same as Step 3 of Algorithm 9.
    S.2.4 Challenges for developing rejection samplers involving Pólya densities
@@
    Appendix S.3 Details of simulation analysis and application study
    S.3.1 Nonparametric binary regression simulation studies
@@
    Table S.3.1: Nonparametric binary regression simulation results comparing computational benefits of different strategies. Entries shown are average over 100 replicates, with Monte Carlo standard error shown in parentheses. ESS, effective sample size of $\lambda$; ESS/sec, ESS divided by running time in seconds; Acc. rate, acceptance rate of $\lambda$.
    Data generation Algorithm settings ESS ESS/sec Acc. rate (%)
    LBP, $\varrho=0.1$ Non-blocked, Indep. M-H Adapted 7.89 (0.35) 0.14 (0.01) 13.60 (0.26)
    Non-adapted 7.13 (0.37) 0.13 (0.01) 12.44 (0.31)
    Blocked, Indep. M-H Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
@@
    Blocked, Particle Gibbs Adapted 358.10 (18.57) 2.33 (0.12) 84.50 (0.42)
@@
    LBP, $\varrho=0.2$ Non-blocked, Indep. M-H Adapted 7.31 (0.32) 0.11 (0.00) 13.42 (0.22)
    Non-adapted 6.58 (0.32) 0.10 (0.00) 12.48 (0.31)
    Blocked, Indep. M-H Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
@@
    Blocked, Particle Gibbs Adapted 360.48 (21.29) 1.96 (0.12) 86.11 (0.40)
@@
    LBP, $\varrho=0.4$ Non-blocked, Indep. M-H Adapted 6.40 (0.29) 0.09 (0.00) 13.01 (0.22)
    Non-adapted 6.56 (0.32) 0.10 (0.00) 12.88 (0.31)
    Blocked, Indep. M-H Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
@@
    Blocked, Particle Gibbs Adapted 468.76 (25.98) 2.52 (0.14) 87.09 (0.29)